# Euclid preparation

### v1

@@
Simulating thousands of Euclid spectroscopic skies
@@
Euclid Collaboration: P. Monaco pierluigi.monaco@inaf.it Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Parimbelli Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Y. Elkhashab Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Salvalaggio Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    T. Castro Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. D. Lepinzan Euclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Sarpa Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Sefusatti Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    L. Stanco Euclid preparationEuclid preparation    L. Tornatore Euclid preparationEuclid preparation    G. E. Addison Euclid preparationEuclid preparation    S. Bruton Euclid preparationEuclid preparation    C. Carbone Euclid preparationEuclid preparation    F. J. Castander Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Carretero Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. de la Torre Euclid preparationEuclid preparation    P. Fosalba Euclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Lavaux Euclid preparationEuclid preparation    S. Lee Euclid preparationEuclid preparation    K. Markovic Euclid preparationEuclid preparation    K. S. McCarthy Euclid preparationEuclid preparationEuclid preparationEuclid preparation    F. Passalacqua Euclid preparationEuclid preparationEuclid preparationEuclid preparation    W. J. Percival Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    I. Risso Euclid preparationEuclid preparationEuclid preparationEuclid preparation    C. Scarlata Euclid preparationEuclid preparation    P. Tallada-Crespí Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Viel Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    Y. Wang Euclid preparationEuclid preparation    B. Altieri Euclid preparationEuclid preparation    S. Andreon Euclid preparationEuclid preparation    N. Auricchio Euclid preparationEuclid preparation    C. Baccigalupi Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Baldi Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Bardelli Euclid preparationEuclid preparation    P. Battaglia Euclid preparationEuclid preparation    F. Bernardeau Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Biviano Euclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Branchini Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Brescia Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Brinchmann Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Camera Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Cañas-Herrera Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    V. Capobianco Euclid preparationEuclid preparation    V. F. Cardone Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Casas Euclid preparationEuclid preparation    M. Castellano Euclid preparationEuclid preparation    G. Castignani Euclid preparationEuclid preparation    S. Cavuoti Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Cimatti Euclid preparationEuclid preparation    C. Colodro-Conde Euclid preparationEuclid preparation    G. Congedo Euclid preparationEuclid preparation    C. J. Conselice Euclid preparationEuclid preparation    L. Conversi Euclid preparationEuclid preparationEuclid preparationEuclid preparation    Y. Copin Euclid preparationEuclid preparation    F. Courbin Euclid preparationEuclid preparationEuclid preparationEuclid preparation    H. M. Courtois Euclid preparationEuclid preparation    A. Da Silva Euclid preparationEuclid preparationEuclid preparationEuclid preparation    H. Degaudenzi Euclid preparationEuclid preparation    G. De Lucia Euclid preparationEuclid preparation    A. M. Di Giorgio Euclid preparationEuclid preparation    F. Dubath Euclid preparationEuclid preparation    F. Ducret Euclid preparationEuclid preparation    C. A. J. Duncan Euclid preparationEuclid preparationEuclid preparationEuclid preparation    X. Dupac Euclid preparationEuclid preparation    S. Dusini Euclid preparationEuclid preparation    A. Ealet Euclid preparationEuclid preparation    S. Escoffier Euclid preparationEuclid preparation    M. Farina Euclid preparationEuclid preparation    R. Farinelli Euclid preparationEuclid preparation    S. Farrens Euclid preparationEuclid preparation    S. Ferriol Euclid preparationEuclid preparation    F. Finelli Euclid preparationEuclid preparationEuclid preparationEuclid preparation    N. Fourmanoit Euclid preparationEuclid preparation    M. Frailis Euclid preparationEuclid preparation    E. Franceschi Euclid preparationEuclid preparation    M. Fumana Euclid preparationEuclid preparation    S. Galeotta Euclid preparationEuclid preparation    K. George Euclid preparationEuclid preparation    B. Gillis Euclid preparationEuclid preparation    C. Giocoli Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Gracia-Carpio Euclid preparationEuclid preparation    A. Grazian Euclid preparationEuclid preparation    F. Grupp Euclid preparationEuclid preparationEuclid preparationEuclid preparation    L. Guzzo Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    S. V. H. Haugan Euclid preparationEuclid preparation    W. Holmes Euclid preparationEuclid preparation    F. Hormuth Euclid preparationEuclid preparation    A. Hornstrup Euclid preparationEuclid preparationEuclid preparationEuclid preparation    K. Jahnke Euclid preparationEuclid preparation    M. Jhabvala Euclid preparationEuclid preparation    B. Joachimi Euclid preparationEuclid preparation    E. Keihänen Euclid preparationEuclid preparation    S. Kermiche Euclid preparationEuclid preparation    B. Kubik Euclid preparationEuclid preparation    M. Kümmel Euclid preparationEuclid preparation    M. Kunz Euclid preparationEuclid preparation    H. Kurki-Suonio Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. M. C. Le Brun Euclid preparationEuclid preparation    S. Ligori Euclid preparationEuclid preparation    P. B. Lilje Euclid preparationEuclid preparation    V. Lindholm Euclid preparationEuclid preparationEuclid preparationEuclid preparation    I. Lloro Euclid preparationEuclid preparation    D. Maino Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Maiorano Euclid preparationEuclid preparation    O. Mansutti Euclid preparationEuclid preparation    O. Marggraf Euclid preparationEuclid preparation    M. Martinelli Euclid preparationEuclid preparationEuclid preparationEuclid preparation    N. Martinet Euclid preparationEuclid preparation    F. Marulli Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    R. Massey Euclid preparationEuclid preparation    E. Medinaceli Euclid preparationEuclid preparation    S. Mei Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Melchior Euclid preparationEuclid preparation    Y. Mellier Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Meneghetti Euclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Merlin Euclid preparationEuclid preparation    G. Meylan Euclid preparationEuclid preparation    A. Mora Euclid preparationEuclid preparation    M. Moresco Euclid preparationEuclid preparationEuclid preparationEuclid preparation    L. Moscardini Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Munari Euclid preparationEuclid preparationEuclid preparationEuclid preparation    R. Nakajima Euclid preparationEuclid preparation    C. Neissner Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S.-M. Niemi Euclid preparationEuclid preparation    C. Padilla Euclid preparationEuclid preparation    S. Paltani Euclid preparationEuclid preparation    F. Pasian Euclid preparationEuclid preparation    K. Pedersen Euclid preparationEuclid preparation    V. Pettorino Euclid preparationEuclid preparation    S. Pires Euclid preparationEuclid preparation    G. Polenta Euclid preparationEuclid preparation    M. Poncet Euclid preparationEuclid preparation    L. A. Popa Euclid preparationEuclid preparation    L. Pozzetti Euclid preparationEuclid preparation    F. Raison Euclid preparationEuclid preparation    A. Renzi Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Rhodes Euclid preparationEuclid preparation    G. Riccio Euclid preparationEuclid preparation    F. Rizzo Euclid preparationEuclid preparation    E. Romelli Euclid preparationEuclid preparation    M. Roncarelli Euclid preparationEuclid preparation    R. Saglia Euclid preparationEuclid preparationEuclid preparationEuclid preparation    Z. Sakr Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    A. G. Sánchez Euclid preparationEuclid preparation    D. Sapone Euclid preparationEuclid preparation    B. Sartoris Euclid preparationEuclid preparationEuclid preparationEuclid preparation    P. Schneider Euclid preparationEuclid preparation    T. Schrabback Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Scodeggio Euclid preparationEuclid preparation    A. Secroun Euclid preparationEuclid preparation    G. Seidel Euclid preparationEuclid preparation    M. Seiffert Euclid preparationEuclid preparation    S. Serrano Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    C. Sirignano Euclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Sirri Euclid preparationEuclid preparation    J. Steinwagner Euclid preparationEuclid preparation    D. Tavagnacco Euclid preparationEuclid preparation    A. N. Taylor Euclid preparationEuclid preparation    I. Tereno Euclid preparationEuclid preparationEuclid preparationEuclid preparation    N. Tessore Euclid preparationEuclid preparation    S. Toft Euclid preparationEuclid preparationEuclid preparationEuclid preparation    R. Toledo-Moreo Euclid preparationEuclid preparation    F. Torradeflot Euclid preparationEuclid preparationEuclid preparationEuclid preparation    I. Tutusaus Euclid preparationEuclid preparation    L. Valenziano Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Valiviita Euclid preparationEuclid preparationEuclid preparationEuclid preparation    T. Vassallo Euclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Verdoes Kleijn Euclid preparationEuclid preparation    A. Veropalumbo Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Weller Euclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Zamorani Euclid preparationEuclid preparation    E. Zucca Euclid preparationEuclid preparation    V. Allevato Euclid preparationEuclid preparation    M. Ballardini Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    C. Burigana Euclid preparationEuclid preparationEuclid preparationEuclid preparation    R. Cabanac Euclid preparationEuclid preparation    M. Calabrese Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Cappi Euclid preparationEuclid preparationEuclid preparationEuclid preparation    D. Di Ferdinando Euclid preparationEuclid preparation    J. A. Escartin Vigo Euclid preparationEuclid preparation    G. Fabbian Euclid preparationEuclid preparationEuclid preparationEuclid preparation    L. Gabarra Euclid preparationEuclid preparation    J. Martín-Fleitas Euclid preparationEuclid preparation    S. Matthew Euclid preparationEuclid preparation    N. Mauri Euclid preparationEuclid preparationEuclid preparationEuclid preparation    R. B. Metcalf Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Pezzotta Euclid preparationEuclid preparation    M. Pöntinen Euclid preparationEuclid preparation    C. Porciani Euclid preparationEuclid preparation    V. Scottez Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Sereno Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Tenti Euclid preparationEuclid preparation    M. Wiesmann Euclid preparationEuclid preparation    Y. Akrami Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Alvi Euclid preparationEuclid preparation    I. T. Andika Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Anselmi Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Archidiacono Euclid preparationEuclid preparationEuclid preparationEuclid preparation    F. Atrio-Barandela Euclid preparationEuclid preparation    S. Avila Euclid preparationEuclid preparation    A. Balaguera-Antolinez Euclid preparationEuclid preparation    P. Bergamini Euclid preparationEuclid preparationEuclid preparationEuclid preparation    D. Bertacca Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Bethermin Euclid preparationEuclid preparation    A. Blanchard Euclid preparationEuclid preparation    L. Blot Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Borgani Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. L. Brown Euclid preparationEuclid preparation    A. Calabro Euclid preparationEuclid preparation    B. Camacho Quevedo Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    F. Caro Euclid preparationEuclid preparation    C. S. Carvalho Euclid preparationEuclid preparation    F. Cogato Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Conseil Euclid preparationEuclid preparation    S. Contarini Euclid preparationEuclid preparation    A. R. Cooray Euclid preparationEuclid preparation    O. Cucciati Euclid preparationEuclid preparation    S. Davini Euclid preparationEuclid preparation    G. Desprez Euclid preparationEuclid preparation    A. Díaz-Sánchez Euclid preparationEuclid preparation    J. J. Diaz Euclid preparationEuclid preparation    S. Di Domizio Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. M. Diego Euclid preparationEuclid preparation    A. Enia Euclid preparationEuclid preparationEuclid preparationEuclid preparation    Y. Fang Euclid preparationEuclid preparation    A. G. Ferrari Euclid preparationEuclid preparation    A. Finoguenov Euclid preparationEuclid preparation    F. Fontanot Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Franco Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    K. Ganga Euclid preparationEuclid preparation    J. García-Bellido Euclid preparationEuclid preparation    T. Gasparetto Euclid preparationEuclid preparation    V. Gautard Euclid preparationEuclid preparation    E. Gaztanaga Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    F. Giacomini Euclid preparationEuclid preparation    F. Gianotti Euclid preparationEuclid preparation    G. Gozaliasl Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Guidi Euclid preparationEuclid preparationEuclid preparationEuclid preparation    C. M. Gutierrez Euclid preparationEuclid preparation    A. Hall Euclid preparationEuclid preparation    S. Hemmati Euclid preparationEuclid preparation    C. Hernández-Monteagudo Euclid preparationEuclid preparationEuclid preparationEuclid preparation    H. Hildebrandt Euclid preparationEuclid preparation    J. Hjorth Euclid preparationEuclid preparation    S. Joudaki Euclid preparationEuclid preparation    J. J. E. Kajava Euclid preparationEuclid preparationEuclid preparationEuclid preparation    Y. Kang Euclid preparationEuclid preparation    V. Kansal Euclid preparationEuclid preparationEuclid preparationEuclid preparation    D. Karagiannis Euclid preparationEuclid preparationEuclid preparationEuclid preparation    K. Kiiveri Euclid preparationEuclid preparation    C. C. Kirkpatrick Euclid preparationEuclid preparation    S. Kruk Euclid preparationEuclid preparation    V. Le Brun Euclid preparationEuclid preparation    J. Le Graet Euclid preparationEuclid preparation    L. Legrand Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Lembo Euclid preparationEuclid preparation    F. Lepori Euclid preparationEuclid preparation    G. Leroy Euclid preparationEuclid preparationEuclid preparationEuclid preparation    G. F. Lesci Euclid preparationEuclid preparationEuclid preparationEuclid preparation    J. Lesgourgues Euclid preparationEuclid preparation    L. Leuzzi Euclid preparationEuclid preparation    T. I. Liaudat Euclid preparationEuclid preparation    J. Macias-Perez Euclid preparationEuclid preparation    G. Maggio Euclid preparationEuclid preparation    M. Magliocchetti Euclid preparationEuclid preparation    C. Mancini Euclid preparationEuclid preparation    F. Mannucci Euclid preparationEuclid preparation    R. Maoli Euclid preparationEuclid preparationEuclid preparationEuclid preparation    C. J. A. P. Martins Euclid preparationEuclid preparationEuclid preparationEuclid preparation    L. Maurin Euclid preparationEuclid preparation    M. Miluzio Euclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Montoro Euclid preparationEuclid preparationEuclid preparationEuclid preparation    C. Moretti Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    G. Morgante Euclid preparationEuclid preparation    S. Nadathur Euclid preparationEuclid preparation    K. Naidoo Euclid preparationEuclid preparation    A. Navarro-Alsina Euclid preparationEuclid preparation    S. Nesseris Euclid preparationEuclid preparation    K. Paterson Euclid preparationEuclid preparation    A. Pisani Euclid preparationEuclid preparation    D. Potter Euclid preparationEuclid preparation    S. Quai Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Radovich Euclid preparationEuclid preparation    G. Rodighiero Euclid preparationEuclid preparationEuclid preparationEuclid preparation    S. Sacquegna Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Sahlén Euclid preparationEuclid preparation    D. B. Sanders Euclid preparationEuclid preparation    D. Sciotti Euclid preparationEuclid preparationEuclid preparationEuclid preparation    E. Sellentin Euclid preparationEuclid preparationEuclid preparationEuclid preparation    L. C. Smith Euclid preparationEuclid preparation    J. G. Sorce Euclid preparationEuclid preparationEuclid preparationEuclid preparation    K. Tanidis Euclid preparationEuclid preparation    C. Tao Euclid preparationEuclid preparation    G. Testera Euclid preparationEuclid preparation    R. Teyssier Euclid preparationEuclid preparation    S. Tosi Euclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparationEuclid preparation    A. Troja Euclid preparationEuclid preparationEuclid preparationEuclid preparation    M. Tucci Euclid preparationEuclid preparation    C. Valieri Euclid preparationEuclid preparation    A. Venhola Euclid preparationEuclid preparation    F. Vernizzi Euclid preparationEuclid preparation    G. Verza Euclid preparationEuclid preparation    P. Vielzeuf Euclid preparationEuclid preparation    N. A. Walton Euclid preparationEuclid preparation
@@
We present two extensive sets of 3500$+$1000 simulations of dark matter haloes on the past light cone, and two corresponding sets of simulated (‘mock’) galaxy catalogues that represent the Euclid spectroscopic sample. The simulations were produced with the latest version of the Pinocchio code, and provide the largest, public set of simulated skies. Mock galaxy catalogues were obtained by populating haloes with galaxies using an halo occupation distribution (HOD) model extracted from the Flagship galaxy catalogue provided by Euclid Collaboration. The Geppetto set of 3500 simulated skies was obtained by tiling a $1.2\,h^{-1}\,\text{Gpc}$ box to cover a light-cone whose sky footprint is a circle of $30^{\circ}$ radius, for an area of 2763 ${\rm deg}^{2}$ and a minimum halo mass of $1.5\times 10^{11}$$h^{-1}\,M_{\odot}$. The relatively small box size makes this set unfit for measuring very large scales. The EuclidLargeBox set consists of 1000 simulations of $3.38\,h^{-1}\,\text{Gpc}$, with the same mass resolution and a footprint that covers half of the sky, excluding the Milky Way zone of avoidance. From this we produced a set of 1000 EuclidLargeMocks on the $30^{\circ}$ radius footprint, whose comoving volume is fully contained in the simulation box. We validated the two sets of catalogues by analysing number densities, power spectra, and 2-point correlation functions, showing that the Flagship spectroscopic catalogue is consistent with being one of the realisations of the simulated sets, although we noticed small deviations limited to the quadrupole at $k>0.2\,h\,\text{Mpc}^{-1}$. We show cosmological parameter inference from these catalogues and demonstrate that using one realisation of EuclidLargeMocks in place of the Flagship mock produces the same posteriors, to within the expected shift given by sample variance. These simulated skies will be used for the galaxy clustering analysis of Euclid’s Data Release 1 (DR1).

### v3

@@
LXXVI. Simulating thousands of Euclid spectroscopic skies
@@
Euclid Collaboration: P. Monaco    G. Parimbelli    M. Y. Elkhashab    J. Salvalaggio    T. Castro    M. D. Lepinzan    E. Sarpa    E. Sefusatti    L. Stanco    L. Tornatore    G. E. Addison    S. Bruton    C. Carbone    F. J. Castander    J. Carretero    S. de la Torre    P. Fosalba    G. Lavaux    S. Lee    K. Markovic    K. S. McCarthy    F. Passalacqua    W. J. Percival    I. Risso    C. Scarlata    P. Tallada-Crespí    M. Viel    Y. Wang    B. Altieri    S. Andreon    N. Auricchio    C. Baccigalupi    M. Baldi    S. Bardelli    P. Battaglia    F. Bernardeau    A. Biviano    E. Branchini    M. Brescia    J. Brinchmann    S. Camera    G. Cañas-Herrera    V. Capobianco    V. F. Cardone    S. Casas    M. Castellano    G. Castignani    S. Cavuoti    A. Cimatti    C. Colodro-Conde    G. Congedo    C. J. Conselice    L. Conversi    Y. Copin    F. Courbin    H. M. Courtois    A. Da Silva    H. Degaudenzi    G. De Lucia    A. M. Di Giorgio    F. Dubath    F. Ducret    C. A. J. Duncan    X. Dupac    S. Dusini    A. Ealet    S. Escoffier    M. Farina    R. Farinelli    S. Farrens    S. Ferriol    F. Finelli    N. Fourmanoit    M. Frailis    E. Franceschi    M. Fumana    S. Galeotta    K. George    B. Gillis    C. Giocoli    J. Gracia-Carpio    A. Grazian    F. Grupp    L. Guzzo    S. V. H. Haugan    W. Holmes    F. Hormuth    A. Hornstrup    K. Jahnke    M. Jhabvala    B. Joachimi    E. Keihänen    S. Kermiche    B. Kubik    M. Kümmel    M. Kunz    H. Kurki-Suonio    A. M. C. Le Brun    S. Ligori    P. B. Lilje    V. Lindholm    I. Lloro    D. Maino    E. Maiorano    O. Mansutti    O. Marggraf    M. Martinelli    N. Martinet    F. Marulli    R. Massey    E. Medinaceli    S. Mei    M. Melchior    Y. Mellier    M. Meneghetti    E. Merlin    G. Meylan    A. Mora    M. Moresco    L. Moscardini    E. Munari    R. Nakajima    C. Neissner    S.-M. Niemi    C. Padilla    S. Paltani    F. Pasian    K. Pedersen    V. Pettorino    S. Pires    G. Polenta    M. Poncet    L. A. Popa    L. Pozzetti    F. Raison    A. Renzi    J. Rhodes    G. Riccio    F. Rizzo    E. Romelli    M. Roncarelli    R. Saglia    Z. Sakr    A. G. Sánchez    D. Sapone    B. Sartoris    P. Schneider    T. Schrabback    M. Scodeggio    A. Secroun    G. Seidel    M. Seiffert    S. Serrano    P. Simon    C. Sirignano    G. Sirri    J. Steinwagner    D. Tavagnacco    A. N. Taylor    I. Tereno    N. Tessore    S. Toft    R. Toledo-Moreo    F. Torradeflot    I. Tutusaus    L. Valenziano    J. Valiviita    T. Vassallo    G. Verdoes Kleijn    A. Veropalumbo    J. Weller    G. Zamorani    E. Zucca    V. Allevato    M. Ballardini    C. Burigana    R. Cabanac    M. Calabrese    A. Cappi    D. Di Ferdinando    J. A. Escartin Vigo    G. Fabbian    L. Gabarra    J. Martín-Fleitas    S. Matthew    N. Mauri    R. B. Metcalf    A. Pezzotta    M. Pöntinen    C. Porciani    V. Scottez    M. Sereno    M. Tenti    M. Wiesmann    Y. Akrami    S. Alvi    I. T. Andika    S. Anselmi    M. Archidiacono    F. Atrio-Barandela    S. Avila    A. Balaguera-Antolinez    P. Bergamini    D. Bertacca    M. Bethermin    A. Blanchard    L. Blot    S. Borgani    M. L. Brown    A. Calabro    B. Camacho Quevedo    F. Caro    C. S. Carvalho    F. Cogato    S. Conseil    S. Contarini    A. R. Cooray    O. Cucciati    S. Davini    G. Desprez    A. Díaz-Sánchez    J. J. Diaz    S. Di Domizio    J. M. Diego    A. Enia    Y. Fang    A. G. Ferrari    A. Finoguenov    F. Fontanot    A. Franco    K. Ganga    J. García-Bellido    T. Gasparetto    V. Gautard    E. Gaztanaga    F. Giacomini    F. Gianotti    G. Gozaliasl    M. Guidi    C. M. Gutierrez    A. Hall    S. Hemmati    C. Hernández-Monteagudo    H. Hildebrandt    J. Hjorth    S. Joudaki    J. J. E. Kajava    Y. Kang    V. Kansal    D. Karagiannis    K. Kiiveri    C. C. Kirkpatrick    S. Kruk    V. Le Brun    J. Le Graet    L. Legrand    M. Lembo    F. Lepori    G. Leroy    G. F. Lesci    J. Lesgourgues    L. Leuzzi    T. I. Liaudat    J. Macias-Perez    G. Maggio    M. Magliocchetti    C. Mancini    F. Mannucci    R. Maoli    C. J. A. P. Martins    L. Maurin    M. Miluzio    A. Montoro    C. Moretti    G. Morgante    S. Nadathur    K. Naidoo    A. Navarro-Alsina    S. Nesseris    K. Paterson    A. Pisani    D. Potter    S. Quai    M. Radovich    G. Rodighiero    S. Sacquegna    M. Sahlén    D. B. Sanders    D. Sciotti    E. Sellentin    L. C. Smith    J. G. Sorce    K. Tanidis    C. Tao    G. Testera    R. Teyssier    S. Tosi    A. Troja    M. Tucci    C. Valieri    A. Venhola    F. Vernizzi    G. Verza    P. Vielzeuf    N. A. Walton
@@
We present two extensive sets of 3500++1000 simulations of dark matter haloes on the past light cone, and two corresponding sets of simulated (‘mock’) galaxy catalogues that represent the Euclid spectroscopic sample. The simulations were produced with the latest version of the Pinocchio code, and provide the largest, public set of simulated skies. Mock galaxy catalogues were obtained by populating haloes with galaxies using an halo occupation distribution (HOD) model extracted from the Flagship galaxy catalogue provided by Euclid Collaboration. The Geppetto set of 3500 simulated skies was obtained by tiling a $1.2\,h^{-1}\,\text{Gpc}$ box to cover a light-cone whose sky footprint is a circle of $30^{\circ}$ radius, for an area of 2763 ${\rm deg}^{2}$ and a minimum halo mass of $1.5\times 10^{11}$$h^{-1}\,M_{\odot}$. The relatively small box size makes this set unfit for measuring very large scales. The EuclidLargeBox set consists of 1000 simulations of $3.38\,h^{-1}\,\text{Gpc}$, with the same mass resolution and a footprint that covers half of the sky, excluding the Milky Way zone of avoidance. From this we produced a set of 1000 EuclidLargeMocks on the $30^{\circ}$ radius footprint, whose comoving volume is fully contained in the simulation box. We validated the two sets of catalogues by analysing number densities, power spectra, and 2-point correlation functions, showing that the Flagship spectroscopic catalogue is consistent with being one of the realisations of the simulated sets, although we noticed small deviations limited to the quadrupole at $k>0.2\,h\,\text{Mpc}^{-1}$. We show cosmological parameter inference from these catalogues and demonstrate that using one realisation of EuclidLargeMocks in place of the Flagship mock produces the same posteriors, to within the expected shift given by sample variance. These simulated skies will be used for the galaxy clustering analysis of Euclid’s Data Release 1 (DR1), and an even larger set of simulations is planned for the next releases.